The hydroxide ion concentration in an aqueous solution at 25°C is 6.1×10-2 M.

The hydronium ion concentration is  M.

The pH of this solution is .

The hydronium ion concentration in an aqueous solution at 25°C is 6.1×10-2M.

The hydroxide ion concentration is  M.

The pH of this solution is .

The pOH is .

The pOH is .

The pH of an aqueous solution at 25°C was found to be 9.40.

The pOH of this solution is .

The hydronium ion concentration is  M.

The hydroxide ion concentration is  M.
Answer #1

[ok] = 6-1X1ö? in - (CH3ot} = 1.64 X10-13h) answer. pH = -log(Hzot] := -loy Clef4x16137 CS Scanned with a (pH = 12.8) answer